1) Given following data on Ghana expenditure in 2003:

Final consumption expenditure by households and non-profit institutions....725.0

Government final consumption expenditure.....231.8

Gross fixed capital formation......175.9

Change in inventories........+4.2

Imports of goods and services........313.2

Net income from abroad.........+22.4

Exports of goods and services.......282.2

Fixed capital consumption (depreciation).......115.3


Calculate:

a. Gross domestic product (at market prices)

b. Net national income


1
Expert's answer
2021-08-30T12:50:02-0400

a) Gross Domestic Product = Consumption + Investment + Government Spending + Net Exports (Exports - Imports) + Change in Inventories

"GDP=C + I + G + X-M+\Delta"I

"GDP=725+231.8+175.9+282.2-313.2+4.2"

"GDP=1,105.90"

b) Net National Income = GDP + Net foreign factor income - depreciation of fixed capital assets

"NNI=GDP+NFFI-D"

"NNI=1,105.90+22.4-115.3"

"NNI= 1,013"